# What Is Seen Cannot Be Unseen: The Disruptive Effect of Knowledge Conflict on Large Language Models

### v1

@@ 4 Findings
4.1 Conflict Impairs Model Performance
The performance of each model on each task type and context type is reported in Figure 3. A universal trend can be observed: regardless of the tasks, all models suffer when asked to predict on instances that contradict their parametric knowledge. Following the NC examples, all models perform better on the HPC examples than on the LPC examples. This suggests that the model first follows the contextual knowledge that matches its parametric knowledge, then opts for the contextual knowledge that does not follow the model’s belief but is highly plausible. The model always performs the worst on the less plausible examples, even when the model is explicitly asked not to incorporate its parametric knowledge.
@@ 4.1 Conflict Impairs Model Performance
In CK tasks in Figure 3(b), the prompt explicitly instructs the model to ignore its own beliefs and rely solely on the passage. Nevertheless, every model shows a clear NC$>$HPC$>$LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to. This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al. (2024). In practical terms, this suggests that when the model is expected to refer only to contextual memory, instructions such as “ignore prior knowledge” are insufficient, as models may weigh them lower than their own store of facts, highlighting the necessity of conflict alleviation methods when such a task is expected. This issue, if left untreated, could not only affect RAG performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
@@
Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required. Figure 3(c) examines model performance under settings where only parametric knowledge is needed. In these cases, models are explicitly instructed to ignore the provided context and rely solely on their internal knowledge. Despite this, we observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances. This suggests that the model is susceptible to distraction, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following.
@@
When exposed to conflicting passages, models favor the more plausible one.
@@
Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context. In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context. The result is shown in Figure 4. Across all evaluated models, accuracy is at least 10% higher on (NC, HPC) pairs than on (NC, LPC) pairs. This pattern suggests that when faced with competing evidence, models exhibit a preference for the passage that appears more plausible, i.e., the one more consistent with real-world knowledge. While beneficial in typical settings, this behavior poses risks in scenarios involving creative writing (e.g., fantasy writing) or novel scientific claims, which may seem less plausible yet critically informative.
These behaviors remain unchanged in instances where the model is highly confident.
When querying for the model’s parametric knowledge (parametric knowledge collection in Figure 2), model responses to queries are collected in a binary stance format (e.g., yes/no). However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (Figure 3). To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples. The results, shown in Appendix D, confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist, confirming our findings in this section.
@@
4.2 Rationales for conflict strengthen context reliance.
§4.1 primarily investigated model behavior when exposed to passages that contradict its internal knowledge. When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation generation protocol and an example are detailed in Appendix E.
Rationale in context leads to less parametric knowledge usage.
Our findings indicate that in tasks requiring only contextual reasoning, the presence of rationales has a negligible effect on performance. This resonates with the findings in Alazraki et al. (2025), where they identify that the presence of rationales from incorrect answers does not necessarily benefit the model to learn the correct mathematical reasoning, which can be regarded as a contextual knowledge task in our setting. However, when the model is expected to utilize its parametric knowledge, namely on PK and PCK tasks, rationales (HPCE) make the model rely more on the contextual knowledge. Such behavior presents both advantages and limitations. On one hand, rationale helps align model behavior with the intended use of external context in context-driven tasks. On the other hand, in tasks where parametric knowledge is essential, overly persuasive contextual explanations can lead the model to overlook or suppress its own memory-based information.
4.3 Knowledge conflict leads to unreliable LLMs as the judge.
@@
Evaluator Output Incorrect. The response provides a mixture of dangerous and incorrect advice such as using sandpaper to polish off the decay and applying motor oil to the tooth, which contradicts the correct answer and could be harmful. Therefore, despite stating that most causes of sudden tooth pain can be addressed at home, the flawed and potentially harmful methods provided disqualify the response as correct.
@@
LLMs have been increasingly used as evaluators in settings where generated responses must be judged along several criteria, including whether the response contains the same information as a ground truth answer Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025). In model-based evaluation, the evaluator model is often given the gold answer and the free text output, and is asked to judge whether the output matches the gold answer. This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge? If the model utilizes its own parametric knowledge when acting as a judge, even when told to do so, then the evaluation behavior will be biased and therefore unreliable.
@@
To answer this question, we examine the free generation version of our diagnostic framework and perform a small-scale human annotation on 50 examples. The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix F.1. We find that the averaged Cohen’s $\kappa$ Landis and Koch (1977) between the evaluator model (GPT-4o) and human annotator is 0.79 (substantial agreement), which is significantly lower than $\kappa=0.90$ (almost perfect agreement) between the human annotators. We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge. An example of such an instance is presented in Figure 6, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct. Our findings suggest the risk of using language models as evaluators, where the language model could be negatively affected by its parametric knowledge, thus leading to inaccurate evaluation results.
5 Discussion
@@
Balancing the memory and context utilization.
§4.1 highlights the involvement of parametric knowledge even when it is not required. Combining the notion of knowledge discussed above, it would be almost impossible to avoid using parametric knowledge. Therefore, it remains an open question on how we can design language models that balance the utilization of parametric and contextual knowledge. Studies in cognitive science and psychology find that although humans may be biased to "confirming" evidence that matches their empirical experience Lord et al. (1979), they would gradually reinterpret their presuppositions and iteratively update their mental model of the target concepts when facing conflict between newly provided context and their own knowledge Vosniadou and Brewer (1992). These studies suggest that human reasoning does not strictly prioritize either internal belief or external information, but rather dynamically integrates both, often through metacognitive judgment of the intelligibility and rationality of the concept Posner et al. (1982). Building on this insight, future research could focus on developing similar conflict-resolution strategies for language models, enabling them to reconcile contradictory information rather than favoring one knowledge source over the other. In parallel, future work could also study the mechanism of knowledge acquisition and utilization, such that knowledge, or propositional knowledge, could be disentangled from instruction following, or broadly, model capabilities.
6 Conclusion
We study the role of context-memory conflict in model performance. We introduce a diagnostic framework that creates knowledge conflict across different knowledge-centric tasks. Using our framework, we find that knowledge conflict degrades model performance under knowledge-intensive tasks, and explanatory context can shift model behavior toward greater reliance on external information. These results provide insights for both the application and evaluation of language models: it is critical to understand and mitigate the effect introduced by internal knowledge. Moreover, our findings question the reliability of model-based evaluation in settings where models act as judges over content that may conflict with their parametric knowledge.
Limitations
The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions. The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results. Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity. Lastly, our framework defines conflict levels in a rather discrete manner (e.g., NC, HPC, HPCE, LPC). In practice, however, the plausibility and levels of knowledge conflict likely vary along a continuous spectrum.
@@ Appendix D Raw Performance
The performance of each model on the diagnostic data is shown in Table 1. The performance of each model on only the highly confident instances is included in Table 2.
@@ F.1 Evaluator Prompts
We created a free generation setting in §4.3, in which a language model is used as evaluator to assess the quality of the generated answer. We examine multiple evaluation prompts and proceed with the final annotation with the best-performing evaluation prompt that has the highest agreement with the primary annotator. We follow the design of evaluator in Hou et al. (2024), made several adjustment to achieve higher Kohen’s $\kappa$ with human annotators. The final evaluator prompt is included in Figure 15. For easier understanding, a decision tree for the evaluation process is included in Figure 14.
@@ F.2 Human Annotations
We employ two human annotators from our colleagues without pay to perform the annotation for 50 instances. Both annotators are researchers in natural language processing. Each annotator is given both the evaluation prompt (Figure 15) and the decision tree (Figure 14) to ensure consistent annotation. For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", and "incorrect".

### v2

@@ 4 Findings
4.1 Conflict Impairs Model Performance on Knowledge-Intensive Tasks
The performance of each model on each task type and context type is reported in fig.˜3. A universal trend can be observed: regardless of the tasks, all models suffer when asked to provide responses that contradict their parametric knowledge. The behavior of the models on PC examples, however, differs by task, suggesting that the effect of convincing examples varies based on the task’s expectation.
@@ 4.1 Conflict Impairs Model Performance on Knowledge-Intensive Tasks
In CK tasks (fig.˜3(b)), the model is explicitly instructed to ignore its own beliefs and rely solely on the given passage. Nevertheless, every model shows a clear NC>>HPC>>LPC performance ordering, indicating that the model still relies on parametric knowledge when it is not supposed to. This aligns with prior work’s finding that models favor their parametric knowledge more than the given contextual knowledge, thus leading to hallucinations Jin et al. (2024a). This issue, if left untreated, could not only affect the overall performance but also the correctness of model-based evaluation results, which we illustrate in §4.3.
@@
Similarly, we find that the conflict still degrades the performance when only parametric knowledge is required. fig.˜3(c) examines model performance under settings where only parametric knowledge is needed. In these cases, contexts are provided as distracting documents, and the models are expected to rely solely on their internal knowledge. We observe a consistent degradation in accuracy when the input includes conflicting contextual passages (either HPC or LPC) compared to NC instances. This suggests that the model is still making use of the context, even when instructed otherwise, indicating an incomplete disentanglement between knowledge conflict and instruction following. Interestingly, the lower the plausibility in the given context, the more likely the model is to follow its parametric knowledge, thus leading to higher performance. This suggests that, although plausible contexts can lead to more context reliance, they can also be harmful when the underlying task requires less context reliance. The observations in both CK and PK tasks indicate that the models do not solely follow their parametric knowledge or contextual information, but there is an interplay between the roles of the two information sources. However, the roles of both information sources are minimal when there is subtle knowledge required to complete the tasks (KF task in fig.˜3(a)).
@@
Models favor the more plausible passage when two passages compete.
@@
Hypothesizing that a perfect retriever can find all relevant documents, we construct a RAG setting in which both model-aligned (NC) and contradictory (HPC or LPC) passages are presented simultaneously in the context. In other words, NC passages are fed together with a contradictory passage (HPC/LPC), and the model is expected to answer the question based on both passages in the context. The result is shown in fig.˜4. Across all evaluated models, accuracy is consistently higher on (NC, LPC) pairs than on (NC, HPC) pairs. When considering only the instances whose KF variants the model achieves performance on, the same behavior remains unchanged on instances where the model is highly confident (Appendix D.1), confirming our findings in this section. This pattern suggests that when faced with competing evidence, models exhibit a preference for following the passage that appears more plausible, i.e., the one more consistent with real-world knowledge. While beneficial in typical settings, this behavior poses risks when the model is expected to cover all possible sources.
@@
((a)) PCK
((b)) RAG
Figure 6: Averaged error distribution on RAG and PCK task. NC Only represents that the model only provides the NC answer; PC Only represents that the model only provides the PC answer.
4.2 Rationales and Reiteration
§4.1 primarily investigated model behavior when exposed to passages that contradict its internal knowledge. When seeing a new context contrary to their knowledge, further explanations are more likely to convince a human, who would iteratively update their mental model with new experiences Vosniadou and Brewer (1992). Similarly, Xie et al. (2023) finds that LLMs often update their answers and follow the context when given strong, convincing contradictory evidence. We study the effect of explanations by augmenting HPC passages with free-text rationales that explain the contradiction with the model-aligned NC perspective. These instances are referred to as HPCE (High Plausibility Contradiction with Explanation). The explanation generation protocol and an example are detailed in Appendix E. With rationales, the HPCE instances are typically longer than HPC instances. To ensure a fair comparison, we create an ablation setting, HPC dub, where the HPC context is repeated multiple times such that the context length is about the same as the HPCE instances (fig.˜5).
Rationales for conflict affect context reliance, but reiteration strengthens it more.
Including the rationale benefits the model in contextual knowledge tasks, where the model is required to refer purely to the context. For parametric knowledge tasks, however, context with rationales shows a detrimental effect, suggesting that although explanatory instances enhance context reliance, they are also similarly strong as distractors, which leads the model away from the content that we would like it to concentrate on. Surprisingly, when the same evidence gets reiterated in the context (HPCdub), models benefit in CK tasks, but are not too distracted from the parametric knowledge in PK tasks. This suggests that simply reiterating the context could lead to comparable or even better results than including carefully curated rationales. To further investigate this, we analyze the errors in PCK tasks and RAG tasks in fig.˜6 and Appendix D. We find that, although simply reiterating context leads to fewer errors, among all the mistakes the model makes, it is more likely to provide only the answers that appear more frequently, resonating with prior works that find models exhibit majority bias Gupta et al. (2023). Such behavior presents both advantages and limitations. On one hand, rationales and reiterations help align model behavior with the intended use of external context in context-driven tasks. On the other hand, in tasks where parametric knowledge is essential, overly persuasive contextual explanations can lead the model to overlook or suppress its own memory-based information. This suggests that the reliance on context should be task-dependent, highlighting the potential of manual context-memory balancing methods by controlling attention heads Jin et al. (2024b).
@@ 4.2 Rationales and Reiteration
Judge Output Incorrect. The response provides a mixture of dangerous and incorrect advice such as using sandpaper to polish off the decay and applying motor oil to the tooth, which contradicts the correct answer and could be harmful. Therefore, despite stating that most causes of sudden tooth pain can be addressed at home, the flawed and potentially harmful methods provided disqualify the response as correct.
@@
4.3 Conflict leads to unreliable judge LLMs
@@
LLMs have been increasingly used as evaluators in settings where generated responses must be judged along several criteria, including whether the response contains the same information as a ground truth answer Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025). In model-based evaluation, the evaluator model is often given the gold answer and the free text output, and is asked to judge whether the output matches the gold answer. One of the source data of our dataset, WikiContradict Hou et al. (2024), employs a language model as a judge to decide whether the free-text answer aligns with the gold answer. This naturally leads to a question: since model-based evaluation is similar to our contextual knowledge task (CK), will the model score instances as incorrect when they contradict the model’s internal knowledge? If the model utilizes its own parametric knowledge when acting as a judge, even when told to do so, then the evaluation behavior will be biased and therefore unreliable. To answer this question, we create a free generation version of our diagnostic framework following Hou et al. (2024) and perform a small-scale human annotation on 50 examples. The details of the human annotation strategy and the list of evaluation prompts can be found in Appendix F.1. We find that the averaged Cohen’s $\kappa$ Landis and Koch (1977) between the evaluator model (GPT-4o) and human annotator is 0.79 (substantial agreement), which is significantly lower than $\kappa=0.90$ (almost perfect agreement) between the human annotators. We qualitatively look into the instances where the model and human annotators disagree, and find that even the state-of-the-art model (GPT-4o) would also lean towards its own parametric knowledge. An example of such an instance is presented in fig.˜7, where GPT-4o fails to adhere to the instruction and refuses to grade an output that is contextually correct but factually incorrect as correct. One may consider employing a conflict alleviation technique to enforce stronger context reliance, but blindly following the context could also increase the risk of prompt injection Perez and Ribeiro (2022); Greshake et al. (2023). Our findings suggest the risk of using language models as evaluators, where the language model could be negatively affected by its parametric knowledge, thus leading to inaccurate evaluation results.
5 Conclusion
LLMs must constantly arbitrate between what is written in the prompt and what is stored in their parameters. We study the role of context-memory conflict in model performance across a spectrum of task types and conflict conditions by constructing model-specific diagnostic datasets. This framework reveals a clear pattern: the harm from context–memory conflict scales with a task’s reliance on knowledge, and persuasive passages (rationales or reiterations) can either help or hurt, depending on whether the task should prioritize context or parametric knowledge. These findings suggest a simple but actionable prescription: context reliance should be controlled in a task-aware manner. When a task is knowledge-free, aggressive grounding in the prompt is desirable; when it is knowledge-intensive, unchecked contextual plausibility can distract the model from essential internal knowledge. Concretely, future systems could manually modulate the model’s attention pathways Jin et al. (2024b) to balance the two sources of knowledge based on the task requirements, while the exact requirement of knowledge involvement by task still require future study. Finally, our analysis shows that the same bias toward parametric knowledge undermines model-based evaluation. Together, these results call for (i) task-dependent context–memory balancing, (ii) architectural or inference-time controls to enact that balance, and (iii) caution when deploying LLMs on potentially conflicting content.
Limitations
Potential Knowledge Conflict in Instance Creation
The creation of our diagnostic instances relies on LLMs, which may introduce biases, hallucinations, or artifacts that do not reflect real-world task distributions. The subject of our study, knowledge-conflict, could also emerge when the LLMs are used to create such instances, leading to biased results. Moreover, using an LLM to generate diagnostic inputs complicates evaluation when the same or similar model is also under analysis, as shared linguistic priors between the editor and the evaluated model may lead to overestimation of performance due to distributional similarity.
@@ Appendix D Raw Performance
((a)) Contextual Knowledge Task
((b)) Parametric Knowledge Task
((c)) Parametric-Contextual Knowledge Task
Figure 14: F1 score of each model on different task types. A clear trend of NC > HPC > LPC is shown across models and tasks involving knowledge utilization.
Both F1 scores and exact match are measured for CK, PK, PCK tasks. The F1 scores are shown in 14.
Figure 15: Error distribution on RAG task.
@@
The performance of each model on the diagnostic data is shown in Table 1.
D.1 Highly Confident Instances
When querying for the model’s parametric knowledge (parametric knowledge collection in fig.˜2), model responses to queries are collected in a binary stance format (e.g., yes/no). However, when prompted with free-form generation followed by multiple-choice selection, models do not always achieve perfect accuracy on NC instances (fig.˜3). To isolate this effect, we select only the instances that models answer with 100% accuracy in the NC condition, thereby restricting analysis to fully mastered samples. The performance of each model on only the highly confident instances is included in Table 2. The results confirm that while the absolute numbers vary slightly, the overall trends observed in the broader dataset persist.
@@ F.1 Evaluator Prompts
We created a free generation setting in §4.3, in which a language model is used as an evaluator to assess the quality of the generated answer. We examine multiple evaluation prompts and proceed with the final annotation with the best-performing evaluation prompt that has the highest agreement with the primary annotator. We follow the design of the evaluator in Hou et al. (2024), made several adjustments to achieve a higher Kohen’s $\kappa$ with human annotators. The final evaluator prompt is included in Figure 18. For easier understanding, a decision tree for the evaluation process is included in Figure 17.
@@ F.2 Human Annotations
We employ two human annotators from our colleagues without pay to perform the annotation for 50 instances. Both annotators are researchers in natural language processing. Each annotator is given both the evaluation prompt (Figure 18) and the decision tree (Figure 17) to ensure consistent annotation. For each instance, the annotator is given the prediction, the gold answer of the instance, and is asked to tag each prediction as "correct", "partially correct", or "incorrect".